OpenStudy (anonymous):

find the range of each function for the given domain. 18- f(x)=2x-7; {-2, -1, 0, 1, 2} 19- g(x)=-4x+1; {-5, -1, 0, 1, 2, 10} 20- h(x)=x^2; {-1.2, 0, 0.2, 1.2, 4} 21- f(x)=8x-3; {-1/2, 1/4, 3/4, 1/8} sorry if this is too much, i just don't really understand how to find the range of a function.. i'd greatly appreciate if you could help me out with the answers. xx

OpenStudy (anonymous):

plug in the domain values in to the function and see what you get it is going to take a while

OpenStudy (anonymous):

\[f(x)=2x-7\] \[f(-2)=2\times (-2)+7=-4+7=3\] so one range value is \(3\) now you have to also find \[f(-1),f(0),f(1),f(2)\] to find the other range values

OpenStudy (anonymous):

Domain relates to x and Range relates to y or f(x)